OpenStudy (anonymous):

the sum of the first 14 terms of the sequnce 1/4 + 1/2 + 1+...

OpenStudy (anonymous):

\[a=\frac{1}{4},r=2\] you get \[\frac{1}{4}\times \frac{2^{15}-1}{2-1}\]

OpenStudy (anonymous):

hold on maybe my exponent is wrong.

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

i dunno i am messing up with the exponent for some reason. we can do this easily \[1+2+4+8+...+2^{11}=2^{12}-1\] and then add \[\frac{3}{4}\] to that

OpenStudy (anonymous):

whats the final answer

OpenStudy (zarkon):

\[2^{12}-1+\frac{3}{4}=\frac{1}{4}\cdot\frac{2^{14}-1}{2-1}\]

OpenStudy (zarkon):

\[=\frac{16383}{4}=4095.75\]

OpenStudy (anonymous):

kay thnks

OpenStudy (anonymous):

i have one more question

OpenStudy (anonymous):

yes i suppose it do, don't it? i messed up the first exponent and wrote 15